# FedCache 2.0: Exploiting the Potential of Distilled Data in Knowledge Cache-driven Federated Learning

### v1

@@ Abstract
Federated Edge Learning (FEL) has emerged as a promising approach for enabling edge devices to collaboratively train machine learning models while preserving data privacy. Despite its advantages, practical FEL deployment faces significant challenges related to device constraints and device-server interactions, necessitating heterogeneous, user-adaptive model training with limited and uncertain communication. In this paper, we introduce FedCache 2.0, a novel personalized FEL architecture that simultaneously addresses these challenges. FedCache 2.0 incorporates the benefits of both dataset distillation and knowledge cache-driven federated learning by storing and organizing distilled data as knowledge in the server-side knowledge cache. Moreover, a device-centric cache sampling strategy is introduced to tailor transferred knowledge for individual devices within controlled communication bandwidth. Extensive experiments on five datasets covering image recognition, audio understanding, and mobile sensor data mining tasks demonstrate that (1) FedCache 2.0 significantly outperforms state-of-the-art methods regardless of model structures, data distributions, and modalities. (2) FedCache 2.0 can train splendid personalized on-device models with at least 28.6 improvement in communication efficiency. Our code is available at https://github.com/poppanda/FedCache2.0.
1 Introduction
Federated Edge Learning (FEL) tak2021federated; duan2023combining represents a specialized form of Federated Learning (FL) yang2019federated designed to operate at the edge of the network, which enables edge devices (clients) to jointly train machine learning models under the coordination of an edge server (server) without sharing raw data. With the growing prevalence of mobile and Internet of Things (IoT) devices coupled with increasing concerns over data privacy, FEL has empowered wide adoption of various on-device Artificial Intelligence (AI) applications, including smart transportation transportation, healthcare healthcare, and recommendation yuan2023federated; guo2021prefer. Despite its promising potential, FEL faces significant challenges in practical deployment, primarily deriving from device-side limitations and the intricate dynamics of device-server interactions in mobile edge networks. Diverse user behaviors and device capabilities necessitate the deployment of scalable and user-adaptive models tailored to heterogeneous hardware specifications and data distributions across devices tak2021federated; duan2023combining; pflsurvey; yu2021toward. These requirements call for FEL to function efficiently across a wide range of device configurations and user preferences. Additionally, communication constraints, such as limited bandwidth and uncertain connections, further complicate these interactions tak2021federated; zhang2022federated; zhu2022resilient. Common devices such as smartwatches, mobile phones, and tablets often struggle to sustain excessive wireless communication or ensure consistent connectivity. These limitations hinder the efficiency of collaborative model updates, thereby impacting the overall performance of the FEL system and the benefits it delivers to users. While a series of works have explored the trade-offs among the aforementioned constraints mills2021multi; wu2023fedict; jin2022personalized; zhu2022resilient; wang2022accelerating, they fall short of addressing all the challenges that are simultaneously prevalent in real-world edge deployments tak2021federated; zhang2022federated.
@@ 1 Introduction
Recently, knowledge cache-driven federated learning (FedCache) wu2024federated has emerged as a game-changer in the FEL field, offering a prevailing paradigm that revolutionizes the mainstream parameters interaction protocol mcmahan2017communication; li2020federated; reddi2021adaptive; spectral; mills2021multi; jin2022personalized; lee2022preservation; wu2024federated and allowing for the joint addressing of the aforementioned challenges. By maintaining a server-side knowledge cache that stores and organizes on-device logits as transferable knowledge, FedCache facilitates communication-efficient and heterogeneous-compatible distillation-based personalized optimization without relying on unrealistic public datasets liu2022communication; wu2022communication. In addition, FedCache eliminates the requirement for multiple devices to remain online simultaneously during training, aligning with the dynamic nature of edge environments. However, the effectiveness of FedCache is limited by the breadth of information available from the remote knowledge cache due to its logits interaction design. Its applicability across various data modalities and application tasks is also restricted by the choice of a task-specific encoder used for extracting sample hashes.
@@
To tackle these limitations, we introduce FedCache 2.0, a novel personalized FEL architecture that improves the performance of heterogeneous on-device models with efficient communication and uncertain connection tolerance. FedCache 2.0 offers a new interaction perspective between devices and the server by incorporating the benefits from both dataset distillation lei2023comprehensive and knowledge cache-driven FL wu2024fedcache, employing privacy-preserving distilled data as the knowledge cached on the server. In this setup, devices perform dataset distillation with the assistance of cached knowledge from the remote server. The distilled data is then shared with the server, ensuring the knowledge cache remains updated with the latest information. To balance system performance and communication efficiency, a device-centric cache sampling strategy is proposed for tailoring transferred knowledge for individual devices within the constraints of available communication bandwidth. The key superiorities of FedCache 2.0 compared with the original FedCache are twofold. First, FedCache 2.0 provides richer information characterization capabilities by storing and transferring distilled synthetic data rather than logits, enabling on-device models to optimize with sufficient server-side information and achieve better precision. Second, FedCache 2.0 adopts a more generalized data anonymization method, enhancing its extensibility to a broader range of data modalities and application tasks. Our proposed architecture maintains the advantages of model heterogeneity allowance, learning personalization, uncertain connection tolerance, and efficient communication from FedCache, while also achieving remarkable performance gains by fully exploiting the knowledge from distilled data.
@@
Contributions. The main contributions of this paper are as follows: (1) We propose FedCache 2.0, a novel knowledge cache-driven federated learning architecture for personalized edge intelligence scenarios, enabling efficient post-distillation data-based knowledge storage, organization, and personalized model training while protecting data privacy. (2) We introduce federated dataset distillation and device-centric cache sampling that matches the FedCache 2.0 architecture. The former generates privacy-preserving and semantically-informed synthetic data, while the latter personalizes knowledge cache sampling to accommodate clients with diverse local data distributions and communication constraints. (3) We conduct comprehensive experiments on five datasets, encompassing image recognition, audio understanding, and mobile sensor data mining tasks. Built upon diversified data heterogeneity, model settings, and application scenarios, FedCache 2.0 not only consistently outperforms state-of-the-art methods (at least 1.7% average User model Accuracy mills2021multi enhancement) in all considered settings, but also achieves better communication efficiency (at least 29.6) compared with baseline algorithms.
2 Related Work
Personalized Federated Learning. A variety of approaches have been developed to tackle the dual challenges of learning and model scale personalization within FL. Differentiated client-side model optimization objectives are implemented in studies such as marfoq2022personalized; shi2023prior; mills2021multi; jin2022personalized; spectral; fedfed, enabling trained models to generalize across clients with varying local data distributions. Novel client-server interaction designs, which depart from the traditional FedAvg mcmahan2017communication, are explored in wu2022communication; wang2023towards; zhu2021data; wu2024exploring; huang2022learn to better accommodate diverse client hardware configurations with differently structured models. Furthermore, hybrid approaches such as wu2024fedcache; wu2023fedict are proposed to simultaneously address both model and learning personalization in FL.
Federated Learning in Edge Computing. The efficiency of executing FL at the network edge has become a hot topic. Research works such as alam2022fedrolex; ijcai2022p399; he2020group; wen2024rtifed investigate the technical frameworks required for running FL algorithms on devices constrained by computational power or storage resources. Device heterogeneity and connection uncertainty in edge environments are tackled by methodologies such as zhu2022resilient; wu2024fedcache; liu2023adaptive; wu2023fedict. Moreover, wang2022accelerating; wu2023agglomerative; deng2023hierarchical; liu2020client; wang2021resource extends FL to a multi-tier architecture involving end-edge-cloud collaborations, enhancing model training efficiency and final performance by leveraging the edge as a bridge between devices and the cloud during the training process.
Federated Learning with Alternative Information. Instead of transmitting model parameters, alternative information is utilized in the FL training process by a series of recent works. Model-agnostic outputs are exchanged between clients and the server in wu2024fedcache; itahara2021distill; huang2022learn; wu2024exploring; li2019fedmd, allowing deployment of customized on-device models across resource-heterogeneous clients. Additionally, methodologies involving uploading mixed or distilled data from clients to servers are proposed in song2023federated; oh2020mix2fld, significantly reducing communication overhead while maintaining client data privacy.
@@ 3 Problem Statement and Reformulation
Background and Preliminary. We consider an FL system deployed at the edge of the network, comprising participating edge devices (clients) coordinated by an edge server (server). Each client owns its local dataset with samples, where each sample are with data dimensions and belong to one of distinct classes. Due to differentiated user behaviors, both the local training an testing datasets among clients are non-independently and identically distributed. Throughout this paper, the terms ’device’ and ’client’ are used interchangeably. Assume that the personalized model parameters of client are denoted as , where indicates the number of parameters in the model of client . Due to system heterogeneity among devices, the required model sizes may vary across clients, such that . Each client has a local objective , which relies on its corresponding local data distribution. The overall goal is to minimize the expected objective across all clients, which is formally expressed as:
Given the instability of device connections in edge environments, multiple clients may not be online simultaneously. Besides, it is essential to minimize the communication overhead between devices and the server under the premise of guaranteeing user model accuracy mills2021multi, saving valuable wireless network resources as well as device energy.
Knowledge Cache-driven Federated Learning. We formulate knowledge cache-driven FL as a distributed optimization problem with the assistance of the remote knowledge cache on the server, that is:
where is the cross-entropy loss, is the softmax function, is the prediction function of the model on client . represents the redundant optimization component of client based on cached knowledge, with corresponding weighting term . As an example, FedCache wang2021knowledge considers model outputs (logits) as knowledge, and optimizes local models based on cached related knowledge, that is:
@@
FedCache 2.0 Optimization Formulation. To address the aforementioned shortcomings of FedCache, FedCache 2.0 is designed to revolutionize the transferred knowledge by shifting from logits to distilled data, as shown in Figure 1 (b). Specifically, local model optimization in FedCache 2.0 is regulated by post-sampled distilled data jointly synthesized by clients, that is:
@@ 4.3 Device-Centric Cache Sampling
FedCache 2.0 | 51.1 | 35.7 | 51.2 | 36.9 | 35.8 | 23.5 |
@@ 5.1 Experimental Setup
Platform. Our experiments are conducted on a high-performance physical server equipped with 12th Gen Intel(R) Core(TM) i7-12700 CPU and multiple NVIDIA GeForce RTX 3090 GPU cards. The server’s memory consists of four 16GB Acer DDR4 modules operating at 2133 MT/s, providing a total of 64GB of RAM. Storage is handled by a KINGSTON SKC3000D2048G solid-state drive.
@@
Datasets. We evaluate the effectiveness of our proposed FedCache 2.0 across various application tasks, including image recognition, audio understanding, and mobile sensor data mining carpineti2018custom. These experiments cover five datasets: which are CIFAR10, CIFAR100 cifar10, CINIC10 cinic10, UrbanSound8K urbansound, and TMD carpineti2018custom. Detailed illustrations of tasks and datasets are provided in Appendix B. Each complete dataset is preprocessed using the distributed data partition strategy from FedML he2020fedml, with a hyper-parameter to adjust the degree of data heterogeneity among clients.
@@
Models. We employ five model structures, considering both deep residual network he2016deep for image data, and fully connected network for numeric data. Detailed information on model structures is illustrated in Appendix C. In addition, we consider both scenarios with homogeneous and heterogeneous on-device models in our experiments, with detailed model settings provided in Appendix D.
@@
Baselines. We compare FedCache 2.0 against the following state-of-the-art methods: MTFL mills2021multi, KNN-Per marfoq2022personalized, spectral co-distillation for personalized FL (SCDPFL) chen2024spectral, FedKD wu2022communication and FedCache wu2024fedcache. These baseline algorithms encompass personalized/multi-task FL methods, FL algorithms addressing dual model heterogeneity and communication efficiency, and FL for edge computing. Implementation details for these algorithms are available in Appendix D.
@@
Criteria. Following mills2021multi, we adopt the average User model Accuracy (UA) as the primary metric for evaluating model precision, focusing on the highest value achieved within 100 communication rounds. In addition, we assess communication efficiency by monitoring the learning curves, measuring average UA against per unit of communication overhead. Detailed elaborations on how communication cost is calculated are provided in Appendix E.
5.2 Performance Evaluation

### v2

# FedCache 2.0: Federated Edge Learning with Knowledge Caching and Dataset Distillation

@@ Abstract
Federated Edge Learning (FEL) has emerged as a promising approach for enabling edge devices to collaboratively train machine learning models while preserving data privacy. Despite its advantages, practical FEL deployment faces significant challenges stemming from device constraints and device-server interactions, necessitating heterogeneous, user-adaptive model training with limited and uncertain communication conditions. Knowledge Cache-driven Federated Learning (FedCache) is a promising architecture that enables communication-efficient and heterogeneous-aware collaborative training in edge computing scenarios. However, previous work is limited by the intrinsic nature of logits-based interactions, leading to performance bottlenecks due to the poor richness of exchanged information for on-device model optimization. To tackle this issue, we introduce FedCache 2.0, a novel personalized FEL architecture that enhances the exchange of optimization insights while delivering state-of-the-art performance with efficient communication. FedCache 2.0 incorporates the benefits of both dataset distillation and knowledge cache-driven federated learning by storing and organizing distilled data as knowledge in the server-side knowledge cache, allowing devices to periodically download and utilize personalized knowledge for local model optimization. Moreover, a device-centric cache sampling strategy is introduced to tailor transferred knowledge for individual devices within controlled communication bandwidth. Extensive experiments on five datasets covering image recognition, audio understanding, and mobile sensor data mining tasks demonstrate that (1) FedCache 2.0 significantly outperforms state-of-the-art methods regardless of model structures, data distributions, and modalities. (2) FedCache 2.0 can train splendid personalized on-device models with at least 28.6 improvement in communication efficiency. Our code is available at https://github.com/poppanda/FedCache2.0.
Index Terms:
Federated learning, edge computing, communication efficiency, dataset distillation, knowledge cache1 Introduction
Federated Edge Learning (FEL) [1, 2] is a specialized form of Federated Learning [3] designed to operate at the edge of the network. It enables edge devices (clients) to jointly train machine learning models under the coordination of an edge server (server) without sharing raw data. With the growing prevalence of mobile and Internet of Things (IoT) devices coupled with increasing concerns over data privacy, FEL has empowered wide adoption of various on-device Artificial Intelligence (AI) applications, including smart transportation [4], healthcare [5], and recommendation [6, 7].
Despite its promising potential, FEL faces significant challenges in practical deployment. One of the primary challenges is the diversity of edge device infrastructures. Devices such as smartwatches, mobile phones, and tablets differ significantly in terms of computational power, storage capacity, and battery life [1, 8]. This diversity calls for the deployment of highly scalable, user-adaptive models that can accommodate heterogeneous hardware specifications across devices [1, 2, 9, 8]. Even among devices with similar hardware configurations, data distributions and user preferences are often highly individualized, further complicating model generalization [10, 11, 12]. Another challenge arises from the communication limitations in mobile edge networks. Edge devices frequently operate under low bandwidth and unreliable network conditions [1, 13, 14], making the transmission of large-scale model parameters impractical due to the precious nature of wireless channels [15, 16]. Moreover, devices are often intermittently online [14], further complicating the coordination required for collaborative model updates. The aforementioned constraints hinder the efficiency of the FEL process, ultimately impacting the systems’ overall performance and the benefits it delivers to users.
Numerous studies have sought to overcome the aforementioned challenges in FEL. Heterogeneous FL [19, 20, 21] focuses on enabling collaborative training across devices with different model architectures, allowing adaptation to varying computational resources and hardware capabilities. Personalized or multi-task FL [11, 17, 22] centers on developing models tailored to individual devices, accommodating diverse user preferences. Communication-efficient FL [16, 23, 24] typically incorporates techniques such as model compression or quantization, reducing the training communication burden with little performance loss. However, each of these approaches addresses only a single challenge, limiting their applicability in real-world edge deployments where multiple challenges coexist.
@@
Given the complexity of the aforementioned challenges, it is evident that a multi-faceted FEL solution is urgently needed. Fortunately, knowledge cache-driven FL (FedCache) [25] offers a prevailing paradigm that revolutionizes the mainstream parameters interaction protocol [26, 27, 28, 29, 11, 22, 30, 31] in FEL. By performing on-device personalized distillation driven by the self-organizing server-side knowledge cache, FedCache facilitates communication-efficient and heterogeneous-compatible collaborative training without requiring multiple devices to remain online simultaneously. However, the performance of FedCache is limited by the logits interaction design, which restricts the amount and quality of information that can be interacted between devices and the server. In addition, the applicability of FedCache across various data modalities and application tasks is also restricted due to its reliance on task-specific encoders.
@@
In this paper, we introduce FedCache 2.0, a novel personalized FEL architecture that improves the performance of heterogeneous on-device models with efficient communication and uncertain connection tolerance, as displayed in Fig. 1. Specifically, FedCache 2.0 revolute the transferred knowledge by shifting from logits to distilled data [32], offering a novel interaction paradigm between devices and the server. In our novel design, devices perform dataset distillation with the assistance of cached knowledge from the remote server. The distilled data is then shared with the server, ensuring the knowledge cache remains updated with the latest information. To balance system performance and communication efficiency, a device-centric cache sampling strategy is proposed for tailoring transferred knowledge for individual devices within the constraints of available communication bandwidth. The key superiorities of FedCache 2.0 compared with the original FedCache are twofold. First, FedCache 2.0 provides richer information characterization capabilities by storing and transferring distilled synthetic data rather than logits, enabling on-device models to optimize with sufficient server-side information and achieve better precision. Second, FedCache 2.0 adopts a more generalized data anonymization method, enhancing its extensibility to a broader range of data modalities and application tasks. Our proposed architecture maintains the advantages of model heterogeneity allowance, learning personalization, uncertain connection tolerance, and efficient communication from FedCache, while also achieving remarkable performance gains by fully exploiting the knowledge from distilled data.
The main contributions of this paper are as follows:
We propose FedCache 2.0, a novel personalized FEL architecture that integrates dataset distillation with knowledge cache-driven federated learning. To our best knowledge, FedCache 2.0 is the first federated learning architecture to achieve state-of-the-art performance while tackling the multifaceted challenges of edge deployment.
We introduce federated dataset distillation and device-centric cache sampling that matches FedCache 2.0 design, facilitating knowledge generation, storage, and organization as well as personalized model training, with data privacy protected.
@@
We conduct comprehensive experiments on five datasets, encompassing image recognition, audio understanding, and mobile sensor data mining tasks. Built upon diversified data heterogeneity, model settings, and application scenarios, FedCache 2.0 not only consistently outperforms state-of-the-art methods (at least 1.7% average User model Accuracy [11] enhancement) in all considered settings, but also achieves better communication efficiency (at least 29.6) compared with baseline algorithms.
@@ 2 Problem Statement and Reformulation
2.1 Background and Preliminary
The main notations of this paper are summarized in Table I. Without loss of generality, we consider an FL system deployed at the edge of the network, comprising participating edge devices (clients) coordinated by an edge server (server). Each client owns its local dataset with samples, where each sample are with data dimensions and belong to one of distinct classes. Due to differentiated user behaviors, both the local training an testing datasets among clients are non-independently and identically distributed. Throughout this paper, the terms ’device’ and ’client’ are used interchangeably. Assume that the personalized model parameters of client are denoted as , where indicates the number of parameters in the model of client . Due to system heterogeneity among devices, the required model sizes may vary across clients, such that . Each client has a local objective , which relies on its corresponding local data distribution. The overall goal is to minimize the expected objective across all clients, which is formally expressed as:
| Notation | Description |
| Number of participating edge devices (clients) | |
| Local dataset of client | |
| The -th sample in the local dataset of client | |
| Number of classes | |
| Personalized model parameters of client | |
| Number of parameters in the model of client | |
| Local objective function of client | |
| Knowledge cache on the server | |
| Cross-entropy loss | |
| Softmax function | |
| Prediction function of the model on client | |
| Redundant optimization component of client based on cached knowledge | |
| Kullback-Leibler divergence loss | |
| The -th knowledge fetched from the knowledge cache for sample index | |
| Distilled data on client | |
| Adaptive sample strategy tailored for client | |
| Feature extractor of the model on client | |
| Classifier of the model on client | |
| Periodically updated random replacement function | |
| Subset of with elements | |
| Dataset distillation objective on client | |
| Local training loss function on client | |
| Label frequency of class on client |
Given the instability of device connections in edge environments, multiple clients may not be online simultaneously. Besides, it is essential to minimize the communication overhead between devices and the server under the premise of guaranteeing user model accuracy [11], saving valuable wireless network resources as well as device energy.
2.2 Knowledge Cache-driven Federated Learning
We formulate knowledge cache-driven FL as a distributed optimization problem with the assistance of the remote knowledge cache on the server, that is:
where is the cross-entropy loss, is the softmax function, is the prediction function of the model on client . represents the redundant optimization component of client based on cached knowledge, with corresponding weighting term . As an example, FedCache [33] considers model outputs (logits) as knowledge, and optimizes local models based on cached related knowledge, that is:
@@ 2.2 Knowledge Cache-driven Federated Learning
| Method |
| MTFL [11] | ✗ | ✓ | ✗ | ✗ | Low | ||||||||||
| kNN-Per [17] | ✗ | ✓ | ✗ | ✗ | Medium | ||||||||||
| SCDPFL [18] | ✗ | ✓ | ✗ | ✗ | High | ||||||||||
| FedKD [16] | ✓ | ✗ | ✓ | ✗ | Medium | ||||||||||
| FedCache [25] | ✓ | ✓ | ✓ | ✓ | High | ||||||||||
| FedCache 2.0 | ✓ | ✓ | ✓ | ✓ | Highest |
2.3 FedCache 2.0 Optimization Formulation
To address the aforementioned shortcomings of FedCache, FedCache 2.0 redesigns the original framework of FedCache by incorporating the benefits from both dataset distillation [32], offering a new interaction paradigm between devices and the server. An overview of the FedCache 2.0 architecture is presented in Figure 2(b), and a detailed comparison with state-of-the-art methods is shown in Table II. Specifically, local model optimization in FedCache 2.0 is regulated by post-sampled distilled data jointly synthesized by clients, that is:
@@ 3.2 Federated Dataset Distillation
Method Source of Code Hyper-Parameters Value MTFL [34] optimizer customized batch size 64 learning rate 0.01 local epoch 1 communicaion round 100 server_lr 0.01 bn_private usyb 0.9 0.999 0.5 kNN-Per [35] optimizer Adam batch size 64 learning rate 0.01 local epoch 1 communicaion round 100 aggregator_type centralized client_type KNNPerClient SCDPFL [36] optimizer Adam batch size 64 learning rate 0.01 local epoch 2 communicaion round 100 0.4 0.3 FedKD [37] optimizer Adam batch size 64 learning rate 0.01 local epoch 1 communicaion round 100 tmax 0.98 tmin 0.95 FedCache [38] optimizer Adam batch size 64 learning rate 0.01 local epoch 1 communicaion round 100 1.5 16 FedCache 2.0 Our Implementation optimizer Adam batch size 64 learning rate 0.01 local epoch 5 communicaion round 15 distillation learning rate 0.001 0.5
@@ 3.3 Device-Centric Cache Sampling
| FedCache 2.0 | 51.1 | 35.7 | 51.2 | 36.9 | 35.8 | 23.5 | |
| =0.5 | Method | CIFAR10/45% | CINIC10/40% | CIFAR100/30% |
| MTFL | N/A | N/A | N/A | |
| kNN-Per | N/A | N/A | N/A | |
| SCDPFL | 17.4G (1.0) | 7.8G (1.0) | 17.6G (1.0) | |
| FedKD | N/A | N/A | N/A | |
| FedCache | N/A | N/A | N/A | |
| FedCache 2.0 | 389.1M (45.8) | 193.8M (41.2) | 609.3M (29.6) | |
| =2.0 | Method | CIFAR10/30% | CINIC10/30% | CIFAR100/15% |
| MTFL | N/A | 39.9G (1.0) | 64.5G (1.0) | |
| kNN-Per | 12.1G (1.6) | N/A | 17.6G (3.7) | |
| SCDPFL | 19.4G (1.0) | 10.6G (3.8) | 15.9G (4.1) | |
| FedKD | N/A | N/A | N/A | |
| FedCache | N/A | N/A | N/A | |
| FedCache 2.0 | 315.9M (62.9) | 473.5M (86.3) | 197.5M (334.4) |
@@ 4.1 Experimental Setup
Platforms. Our experiments are conducted on a high-performance physical server equipped with 12th Gen Intel(R) Core(TM) i7-12700 CPU and multiple NVIDIA GeForce RTX 3090 GPU cards. The server’s memory consists of four 16GB Acer DDR4 modules operating at 2133 MT/s, providing a total of 64GB of RAM. Storage is handled by a KINGSTON SKC3000D2048G solid-state drive.
@@
Datasets. We evaluate the effectiveness of our proposed FedCache 2.0 across various application tasks, including image recognition, audio understanding, and mobile sensor data mining [39]. These experiments cover five datasets: which are CIFAR10, CIFAR100 [40], CINIC10 [41], UrbanSound8K [42], and TMD [39]. Detailed illustrations of tasks and datasets are provided in Appendix B. Each complete dataset is preprocessed using the distributed data partition strategy from FedML [43], with a hyper-parameter to adjust the degree of data heterogeneity among clients.
@@
Models. We employ five model structures, considering both deep residual network [44] for image data, and fully connected network for numeric data. In addition, we consider both scenarios with homogeneous and heterogeneous on-device models in our experiments, with detailed model settings provided in section 4.2.
@@
Baselines. We compare FedCache 2.0 against the following state-of-the-art methods: MTFL [11], KNN-Per [17], spectral co-distillation for personalized FL (SCDPFL) [18], FedKD [16] and FedCache [25]. These baseline algorithms encompass personalized/multi-task FL methods, FL algorithms addressing dual model heterogeneity and communication efficiency, and FL for edge computing.
@@
Criteria. Following [11], we adopt the average User model Accuracy (UA) as the primary metric for evaluating model precision, focusing on the highest value achieved within 100 communication rounds. In addition, we assess communication efficiency by monitoring the learning curves, measuring average UA against per unit of communication overhead. Detailed elaborations on how communication cost is calculated are provided in Appendix D.
4.2 Implementation Details
Data Partition. For all experiments conducted in this paper, we utilize the data partition strategy provided by FedML [43] to simulate decentralized datasets for clients. We adopt the hyper-parameter to control the degree of data heterogeneity. A smaller number of results in a greater degree of heterogeneity among clients. Notably, the training and testing datasets for the same client have identical distributions, whereas different clients typically have different data distributions, as illustrated in Figure 4.
Model Configurations. We conduct experiments in both model-homogeneous and model-heterogeneous settings respectively, with detailed information on model structures provided in Appendix C. For image recognition experiments with homogeneous models, all clients employ ResNet-L. For image recognition experiments with heterogeneous models, models on clients are evenly distributed among ResNet-S, ResNet-M, and ResNet-L. Note that for this set of experiments, we only consider FedKD [16] and FedCache [25] as baseline algorithms to compare with. This choice is based on the fact that MTFL [11], kNN-Per [17], and SCDPFL [18] support only homogeneous models across clients. In addition, ResNet-T is adopted to facilitate the interaction of model parameters between clients and the server within the FedKD algorithm in all image recognition experiments. For audio understanding on UrbanSound8K [42] and mobile sensor data mining on TMD [39], we exclusively employ the model-homogeneous setting, with all clients deploying a uniform FCN model, whose structures are denoted as FCN-U and FCN-T, respectively.
Hyper-parameter Settings. We provide a default setting in Table III, which is implemented in all experiments unless stated otherwise. Note that the table only showcases a subset of hyperparameters for baseline algorithms; those not mentioned retain their original settings as specified in the corresponding open-source code. Exclusively, we uniformly set the learning rate to 0.1 and the communication round of kNN-Per to 200 for experiments on mobile sensor data mining.
4.3 Evaluation on Image Classification